OpenStudy (gizmo):

X^3+512=0 real or imaginary solution by factoring

OpenStudy (anonymous):

\[512 = 2^9 = 4^3\] so you have \[x^3+4^3\] which can be factored as \[(x+4)(x^2-4x+16)\] can you solve it from here?

OpenStudy (anonymous):

real